OpenStudy (unbelievabledreams):

Neon-20 Oxygen-19 Sodium-22 Nitrogen-16

OpenStudy (chmvijay):

can u tell atomic no of all atoms u mentioned

OpenStudy (unbelievabledreams):

Neon = 10 Oxygen = 8 Sodium = 16 Nitrogen = 7 Fluorine = 9

OpenStudy (anonymous):

Fluorine: 19-9=10 Neon: 20-10=10 Oxygen: 19-8=11 Sodium: 22-11=11 Nitrogen: 16-7= 9 So Fluorine-19 has the same amount of neutrons as Neon-20